Minutes of the 39^{th} Annual General Meeting of the Dublin University Rifle Club

Moderator John Keeney Minutes taker Mark Dennehy Present John Keeney (Captain),

Alan Smyth (Secretary), Benjamin Jones (Treasurer), Mark Dennehy (Chief Range Officer/Match Secretary), Ray Reilly, Keith White, Michael Madigan, Joseph Lalor, Adrian Costigan, Peter Byrne, Conor McDermottroe, Una Vejsbjerg, Deirdre Ninaber, Steven Watterson, Evelyn Timothy , Willie O’Connor, Shane Kelly, David Gallagher, Liam Murray (late arrival), Meredith Lulling (late arrival), Gerard Ryan (late arrival)

Absent(with apologies) Jason Sibley, Frank De Salle, Kenneth Suzuki

Absent Terry McAuley, Cathy Doyle, Geraldine McAuley Dr. Trevor West

Date 9,May 2001 1800 – 2030 Distribution Terry McAuley (DURC President and Director of Sport),

Dr. Trevor West (DUCAC Chairman), Cathy Doyle (DUCAC Administrative Officer), Geraldine McAuley (DUCAC Administrative Officer), Michelle Bennett (DUCAC Sports Facilities Committee Chairperson),

Helen Murphy (DUCAC Sports Facilities Committee Secretary),

Overview of topics

1 Minutes of the previous AGM 3

2 Captains Report 3

3 Treasurers Report 4

4 Chief Range Officers Report 4

5 Match Secretarys Report 5

6 Constitutional Changes 6 6.1 Article 5 amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 6.2 Article 8 amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 6.3 Article 9 amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.4 Article 12 amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.5 Article 15a amendment . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.6 Article 15c amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 6.7 Article 16 amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.8 Article 18 amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.9 Article 19 amendment . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.10 Article 19b amendment . . . . . . . . . . . . . . . . . . . . . . . . . . 8 6.11 Article 20a amendment . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.12 Article 20b and Article 21 amendments . . . . . . . . . . . . . . . . . . 9

7 Appointment of Vice-Presidents 9

8 Election of next years Club Officers and Committee 9 8.1 Captain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 8.2 Secretary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 8.3 Treasurer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 8.4 Ordinary Committee Members . . . . . . . . . . . . . . . . . . . . . . . 10

9 A.O.B. 12

10 Prizegiving Ceremony 12

1 Minutes of the previous AGM

1 Minutes of the previous AGM

Alan read the minutes of the previous AGM. Much laughter ensued when it transpired that the minutes were incomplete and also that some things proposed in the previous AGM (bringing our fullbore rifles to Northern Ireland from Bisley for example) had been woefully underestimated in complexity. More hilarity when John Keeney’s name was misspelt three times in the minutes.

AGM Minutes were proposed by John Keeney, seconded by Willie O’Connor (who took the minutes at the previous AGM), and were uncontested.

Matters Arising

Ray Reilly enquired as to the status of the transfer of the fullbore rifles in Bisley to Northern Ireland. The consensus was that we may do so in future but due to the complexities of the procedure and workload on Committee this year, it had not been done. Some new members were surprised to learn we even had such rifles ...

The Bisley trip was mentioned, with John Keeney, Ben Jones and Alan Smyth named as going; Mark Dennehy as a provision if finances permitted (he was attending the World Cup in Munich in June as Team Manager).

2 Captains Report

The Captains Report (see attached report) was read by John Keeney; comments were made on the success of the club this year, and how it was a recovery from a few poor years for the club. It was mentioned that the club has, for the first time in its history, both the 50m and 25yd prone rifle national champions in its membership. Winning of colours was also mentioned, as was our Air Rifle Open, the shooting of three maximum possible scores in prone rifle, and the delivery of five new shooting jackets.

On the downside, mention was made of an incident with one of the clubs members which lead to his suspension from the club; of the continuing difficulties with the new sports complex, specifically the way in which the clubs requests for involvement in the planning stages of the new range have been ignored; and the loss of some committee members due to other commitments and illness; a specific mention of Ms. Sharon Wall here, and our wishes for her speedy recovery were heartily seconded by all present.

The Foot and Mouth crisis and its effects on the 50m prone calendar was discussed - the DURC 50m Open will go ahead, but quite late.

Thanks were given to members of the Committee (and Mark Dennehys driving skills suitably mocked). Credit was given where due and John thanked the Range Officers, Cathy and Geraldine from DUCAC, and Security.

Mention was made of how our next year was to be our 40^{th} anniversary and that we had plans underway to celebrate this; Also of the discovery of a letter in the library archives which dated competitive target shooting in the college to 1842. Proposed : John Keeney Seconded : Ben Jones

3 Treasurers Report

Accepted without contest

3 Treasurers Report

The Treasurers Report (see attached Report) was read by Ben Jones. In short, we are broke, but not as badly as last year. The details of the accounts were read through by Ben; along the way managing to dig deep holes over Alans jacket, and provoking a recollection of the Great UCD Sandwich Saga.

The report was interrupted by the lecture hall turning off the lights automatically because everyone was being quite still, so a period of calisthenics was held.

Notes on bank charges for conversion to Euros were brought up by Willie O’Connor; the fact that we have to pay these was seen as strange until Ben pointed out that we can’t use electronic transfer of funds due to the financial structure of DUCAC. Hopefully the introduction of the Euro as sole legal tender next year will reduce this cost.

Matters Arising

Peter Byrne asked about the costs of hiring Luce Hall for the Air Open as opposed to UCD: The consensus was that UCD gave one large bill while Luce Hall incurred many smaller ones to give a larger total and more administrative headaches. To continue to use UCD as the venue for the Air Open was accepted as the best available solution.

Ray Reilly queried the subsidising of the entry fees for the summer competitive season: Ben replied that we could only fully cover one or two competitions and from then on a partial subsidy was all we could afford.

A general query as to why the finances were so low was answered by Ben as being the result of last years grant not being fully spent (resulting in lower grants this year) coupled with a record rise in membership, stretching the clubs budget to the limit.

Ray Reilly suggested approaching the Trinity Trust for funding : Ben replied that we had and they didn’t want to restart funding for sports clubs, that they felt that was DUCACs domain; Una Vejsbjerg pointed out that they preferred to fund Capital investments as opposed to operational costs; Mark Dennehy pointed out that we were in the process of approaching outside companies for sponsorship.

Keith White queried whether future bills between the AGM and the next DUCAC grant were covered : Ben relied that they were and details of that were in his report. Proposed : Ben Jones Seconded : Mark Dennehy Accepted without contest

4 Chief Range Officers Report

At this point, Ben Jones takes over the minuting of the AGM so Mark Dennehy can read the CROs report and the Match Secretarys Report.

This CRO report (see attached Report) marks the return of the CRO report after an absence of almost a decade, prompted by a recent foray into the old Club records by Mark Dennehy.

Liam Murray, Meridith Lulling arrive

5 Match Secretarys Report

Notes were made on membership numbers and demographics; Activity levels and the standard of shooting as judged by the Ladder averages; A graph of membership since 1975 was presented and explained; Ammunition usage levels and breakdowns into Training and Competition were stated; Congratulations on excellent performances was given to Ray Reilly, John Keeney, Mark Dennehy. Gerald Ryan arrives

The RO and Coaching courses introduced by the NTSA and NRPAI were explained; the club’s wish to see all its ROs nationally certified before the next AGM was stated.

ROs who had left were mentioned and those who had volunteered this year. RO drive brought our RO numbers back to an acceptable level. Specific thanks extended to Keith White for DIY on range, Air Open statistics and his programs (KADA and SCORP); to the “Transport Officers” Ray Reilly, Mark Dennehy, Jason Sibley, Alan Smyth and Peter Byrne; and to Keith White, Ben Jones and Mark Dennehy for cleaning the Range at the start of the year.

Official thanks from the CRO were also extended to the Chief Steward; the Director of Buildings and his office; the Security Guards; the Gardai, specifically Dan Quill and John Rafter; to the Firearms departments of Northern Ireland and the Isle of Man, specifically Hillary Orr and Pam Temple; to DUCAC and the Sports Department, specifically Cathy and Geraldine; to Rosemary Welsh of the CS department for her sufferance of Mark and John using college equipment at times in the name of the Club; and to the club members who shot during the year. Proposed : Mark Dennehy Seconded : John Keeney Accepted without contest

5 Match Secretarys Report

The full list of results is in the attached report. Mark summarised (insofar as was possible) the results for the year (it was a lengthy report because of a successful year). 24 competitions attended, 4 at International level. 9 further internal competitions. Hosted only one open due to FMD - the second to be held later this year.

Items of note were Ray Reilly being the National 50m Champion, setting a club record score of 582 and his performance at several International shoots including two World Cups; Steven Watterson’s 562 on a borrowed rifle at his first Irish Air Rifle shoot; Michael Madigan who won the Rathdrum Rifle and Pistol Club Turkey shoot; the team who shot in the Clay Pigeon Colours against UCD despite confusion over the final result, especially Peter Byrne who shot a near-perfect score; Mark Dennehy for winning the National 25yd Championships; and Conor McDermottroe for being promoted to class C after the minimum 3 competitions, which is an exceptional achievement.

It was noted that the Club has never had (in the past 20 years of records at least) the National 25yd Champion in its ranks, and that this year we had both the 25yd and the 50m champions as DURC members.

6 Constitutional Changes

Thanks were also given to Steven Watterson as an Isle of Man representative for the enormous assistance that members of clubs had given when Mark went there for the Easter shoot, where he set a personal best score of 574, winning his class.

Postal competitions were mentioned, but specifically that as of next year, to prevent people registering for competitions and not shooting them, sanctions would be imposed; namely that if the cards for the competition were not shot on time and the shooter disqualified as a result, he or she would have to repay the club their entry fee.

The final scores for the Trinity Top Ten were officially announced - commiserations to Ray Reilly, who was eleventh, but it was noted that he was second last year.

Due to the length of the report, it was decided to postpone the reading of the internal competition results until the prizegiving ceremony. Proposed : Mark Dennehy Seconded : John Keeney Accepted without contest Joe Lalor left with

apologies

6 Constitutional Changes

The current AGM being unconstitutional due to its ratification at a special meeting held in October (Article 21 Paragraph b required all General Meetings to be held between November and May), the first issue was to officially adopt the current constitution. Proposed : John Keeney Seconded : Mark Dennehy All in favour

Here began the longest session of the meeting, which was frequently complained about by Gerald Ryan and Liam Murray. Gerald Ryan proposed at one point that all amendments be voted on en masse; the motion was not seconded.

Each proposed amendment (the original Constitution, the proposed amendments and the resulting Constitution are attached) was then voted on in turn. All amendments were proposed by John Keeney and seconded by Mark Dennehy. All amendments were initially read by Ben Jones with occasional comment from other officers and then the floor was opened for comment: a vote was then called for by a show of hands. A headcount was car-

ried out at this point; 20 members being in attendance, a majority of 15 was required to pass an amendment, as per Article 19 paragraph b of the constitution

6.1 Article 5 amendment

For  Against  Abstention
18  2  0

Motion carried

6.2 Article 8 amendment

For  Against  Abstention
20  0  0

Motion carried

6.3 Article 9 amendment

For  Against  Abstention
20  0  0

Motion carried

6.4 Article 12 amendment

Ray Reilly asked about the constitution’s text about the Selection Committee. It was agreed that it was important that the clubs internal committees and structure be documented, but not in the Constitution.

For  Against  Abstention
20  0  0

Motion carried

6.5 Article 15a amendment

For  Against  Abstention
20  0  0

Motion carried

6.6 Article 15c amendment

Keith White asked if the Chief Steward had been notified of this proposed change; it was agreed that such notification was not necessary as it was an internal club matter until such time as the committee wished or required his assistance.

Ray Reilly asked if the Gardai had been notified; the answer was the same.

Willie O’Connor stated he was against the modification as he felt the article referred to security threats and not a simple problem such as a drunk member arriving at the club door, and that in such cases mandatory involvement of College and other authorities was necessary; this was an important point but the committee noted that in such cases, outside authorities would be called in any event; and that the consequences to a student who fell afoul of this amendment through honest mistake (ie turning up on the range after a single drink) would be punished too severely with a permanent mark on their academic and legal records.

However the point was serious enough that it was agreed to abandon the amendment and to resubmit it at a later meeting after some reworking.

For  Against  Abstention
0  20  0

Motion rejected

6.7 Article 16 amendment

This motion was abandoned due to an error when the amendment was drawn up.

For  Against  Abstention
1  19  0

Motion rejected

6.8 Article 18 amendment

Keith White objected to this motion, as well as several others, on the grounds that the amendments had been prepared over only three or four days. It was later agreed that he would lead a focus group to examine the Constitution, which is poorly written in some areas, and that this group would propose its amendments at the next General Meeting.

For  Against  Abstention
17  1  2

Motion carried

6.9 Article 19 amendment

Comments were made that the amendment allowed the possibility of nonsensical disciplinary action; it was pointed out by Mark Dennehy that the purpose of the constitution here was to define mechanism and not policy.

More serious discussion was on whether or not the wording of the amendment was adequate; the level of flexibility given to committee was difficult to regulate accurately. After considerable discussion, it was agreed that the motion would remain tabled, but that a further amendment to this article would be presented at the next General Meeting; the need for the amendment being stated as very great by the committee who had had occasion that year to exercise this article.

For  Against  Abstention
17  2  1

Motion carried

6.10 Article 19b amendment

For  Against  Abstention
20  0  0

Motion carried

7 Appointment of Vice-Presidents

6.11 Article 20a amendment

This amendment was withdrawn by committee for further reworking and resubmittal at the next General Meeting.

6.12 Article 20b and Article 21 amendments

These were proposed and voted on as a single block with no objections save Keith’s previously stated ones.

For  Against  Abstention
18  2  0

Motion carried Many sighs of relief once this voting was concluded!

7 Appointment of Vice-Presidents

Mr. Ray Reilly and Mr. Keith White were invited by John Keeney (after some fumbling with the appropriate wording) to become Vice-Presidents of the Club. This being deemed by all to be one of the highest honours the Club can bestow, offered in these cases for long service and long-standing standards of excellence displayed by both shooters. Both accepted.

John thanked both and then noted that they were now to continue to “lead the club from behind” (which prompted much mirth from Messrs Murray and Ryan). Ray Reilly was to be appointed to inventory and curate the Club’s collection of trophies and Keith White was to lead a focus group to reexamine the Constitution to reword it more legibly and clearly.

Both shooters received applause from those present.

8 Election of next years Club Officers and Committee

Willie O’Connor pointed out that all Club Officers had to be registered students according to DUCAC rules, but that this did not apply to Ordinary Committee Members.

8.1 Captain

Nominees :

Alan Smyth Proposed : John Keeney Seconded : Liam Murray

8 Election of next years Club Officers and Committee

Steven Watterson Proposed : Ger Ryan Seconded : Willie O’Connor

The building lights again went out - the resulting arm-waving gave the impression of the greatest interest in the Captaincy of the club in many years ...

Voting was then carried out by written secret ballot.

Name  Votes
Alan Smyth  7
Steven Watterson  11
Abstentions  2

Steven was deemed elected as Captain for the 2001-2002 year.

8.2 Secretary

Nominees :

Mark Dennehy Proposed : Ben Jones Seconded : John Keeney

There being no other nominees, Mark was deemed elected - then Liam Murray asked if this was correct procedure, so a show of hands was asked for and Mark was unanimously accepted.

8.3 Treasurer

Nominees :

Ben Jones Proposed : John Keeney Seconded : Conor McDermottroe

There being no other nominees, a show of hands deemed Ben unanimously elected as Treasurer.

8.4 Ordinary Committee Members

Nominees :

Alan Smyth Proposed : John Keeney Seconded : Liam Murray

Liam Murray Proposed : Ger Ryan Seconded : Adrian Costigan

Deirdre Ninaber Proposed : Mark Dennehy Seconded : Alan Smyth

Conor McDermottroe Proposed : Steven Watterson Seconded : John Keeney

John Keeney Proposed : Ben Jones Seconded : Mark Dennehy

Peter Byrne Proposed : Liam Murray Seconded : Conor McDermottroe

Shane Kelly Proposed : Steven Watterson Seconded : Keith White

Jason Sibley Proposed : Alan Smyth Seconded : John Keeney

At this point, John spoke on the level of work and commitment that being a Committee Member required and asked that each nominee consider this in light of their academic work: each was then polled in turn to see if they accepted the nomination. At this point, Peter Byrne and Shane Kelly withdrew their nomination.

A written secret ballot was then held :

Name  Votes
Alan Smyth  16
Liam Murray  10
Deirdre Ninaber  16
Conor McDermottroe  15
John Keeney  16
Jason Sibley  10

9 A.O.B.

At this point, Alan Smyth, Deirdre Ninaber, Conor McDermottroe and John Keeney were deemed elected; a further secret written ballot was then held between Jason Sibley and Liam Murray :

Name  Votes
Liam Murray  8
Jason Sibley  10

Jason was deemed to be the fifth committee member. At this point, Mr Ryan

indicated his unhappiness at the length of the AGM

The Committee for the next (and 40^{th}) year of the Club was announced as :

Captain  Steven Watterson
Secretary  Mark Dennehy
Treasurer  Ben Jones
Other Committee Members  Deirdre Ninaber

Conor McDermottroe John Keeney Alan Smyth Jason Sibley

9 A.O.B.

An inquiry was made regarding the presentation of the prizes which were on display on the table; it was announced that they would be given out back on the range during the reception.

At this point, the AGM was adjourned to the range for a reception and prizegiving.

10 Prizegiving Ceremony

Various prizes were given out to the winners of the internal competitions (see the attached Match Secretarys report).

The outstanding newcomer award was given, after a lengthy debate, to Conor Mc-Dermottroe for his outstanding Target Rifle performance in competition and general commitment to the Club. However it was noted that the decision had been a very difficult one: Liam Murray and Adrian Costigan had given great competition for this prize and were specifically mentioned. Unfortunately the actual silver plate for the award was still in the possession of Miss Anna Hickey, who was in Queens at the time: it was agreed that the plate would be retrieved and engraved for Conor.

Colours were awarded to Mark Dennehy (who was suitably chuffed and found it difficult to stop wearing the scarf for the remainder of the night) for his performance in Air and Free rifle through the year and for his commitment to the club’s operation during the year.

10 Prizegiving Ceremony

Mark said that this year had had more excellent beginners than any other year that he had seen.

The reception then began. This concluded the prizegiving and the Annual General Meeting.